Given f(x) = –x3 – 3x2 + 3x + 7, find f(–1), f(0), and f(2)

Answer by checkley71(8403)

F(X)=-X^3-3X^2+3X+7
F(-1)=-(-1)^3-3(-1)^2+3*-1+7
F(-1)=-(-1)-3(1)-3+7
F(-1)=1-3-3+7
F(-1)=8-6
F(-1)=2 ANSWER
-------------------------------------
F(0)=-0^3-3*0^2+3*0+7
F(0)=0+0+0+7
F(0)=7 ANSWER
-------------------------------------
F(2)=-(2)^3-3*2^2+3*2+7
F(2)=-8-3*4+6+7
F92)=-8-12+6+7
F(2)=-20+13
F92)==-7 ANSWER.